OpenStudy (anonymous):

logarithmic equation, solve for x: 43 = 11 ln (4x+5) (my work in reply)

OpenStudy (btaylor):

since 43/11 = ln(4x+5), e^(43/11) = 4x+5 so, 49.9=4x+5 44.9=4x x=11.2

OpenStudy (anonymous):

My answer isn't correct. What am I doing wrong? \[43 = 11 \ln(4x+5)\] \[e^{43} = 11(4x+5)\] \[e^{43} = 44x + 55\] \[e^{43} - 55 = 44x\] \[x = \frac{ e^{43}-55 }{ 44 }\]

OpenStudy (btaylor):

you need to divide by 11 first.

OpenStudy (anonymous):

Answered before I finished typing out my work, haha. That was it. ok, so I need to remember to get the log by itself first? I got: \[x=\frac{ e^{43/11} - 5 }{ 4 } \approx 11.2\] Thank you!

OpenStudy (btaylor):

yup. good job